OpenStudy (anonymous):

Please help me!!!!! Determine a mathematical model that represents this data by examining the patterns in the table. Number of Folds Number of Sections 0 0 1 2 2 4 3 8 4 16 5 32 6 64

OpenStudy (helder_edwin):

if F=number of folds and S=number of sections then for F>0 \[ \large S=2=2^1\qquad F=1 \] \[ \large S=4=2^2\qquad F=2 \] \[ \large S=8=2^3\qquad F=3 \] \[ \large S=16=2^4\qquad F=4 \] \[ \large S=32=2^5\qquad F=5 \] \[ \large S=64=2^6\qquad F=6 \]